1.4.23 Question Help A salesperson obtained a systematic sample of size 25 from a list of 500 clients. To do so, he randomly selected a number from 1 to 20, obtaining the number 13. He included in the sample the 13th client on the list and every 20th client thereafter. List the numbers that correspond to the 25 clients selected. Choose the correct answer below OA. 25, 45, 500 O B. 20, 33,.., 332 ( C. 13, 33, , 488 D. 13, 33, ..., 493
Answer #1

initial number = a =13

Increase by = d = 20

This is a arithmetic progression with n=25.

End number = a+(n-1)d=13+(25-1)*20=493

Ans : Option D.